OpenStudy (anonymous):

Simplify: [x^(3/2)] [(x+x^(5/2) - x^2 )]

OpenStudy (anonymous):

\[\left(x^{\frac{3}{2}}\right)\left(x+x^{\frac{5}{2}}-x^2\right)\] this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Distribute ; remember that \[x^a * x^b= x^{a+b}\]

OpenStudy (anonymous):

\[x^{\frac{3}{2}}*x +x^{\frac{3}{2}}*x^{\frac{5}{2}}-x^{\frac{3}{2}}*x^2\]

OpenStudy (anonymous):

is taht the end of the problem? teh most simplifed it can get.

OpenStudy (anonymous):

no, \[x^{\left(\frac{3}{2}+1\right)}+x^{\left(\frac{3}{2}+\frac{5}{2}\right)}-x^{\left(\frac{3}{2}+2\right)}\]

OpenStudy (anonymous):

i see so is teh answer: x^(5/2) = x^4 - x^ (7/2)